OpenStudy (freemap):

3. When a boat traveled downstream from Town A to Town B, the trip took 3 h. When the same boat traveled upstream from Town B to Town A, the trip took 3.6 h. For each trip, the speed of the boat and the water current were unchanged. Let x represent the speed of the boat and let y represent the speed of the water.

OpenStudy (freemap):

(a) Write an expression for the distance traveled downstream using 3 h for the time. Then write an expression for the distance traveled upstream using 3.6 h for the time. (b) Set the expressions in Part (a) equal to each other. Then solve the equation for y. Show your work. (c) What percent of the boat’s speed is the water current?

OpenStudy (phi):

have you learned rate * time = distance or speed * time = distance or velocity * time = distance ?

OpenStudy (freemap):

yes I have

OpenStudy (phi):

For traveled downstream first, what is the speed of the boat going with the current?

OpenStudy (freemap):

I don't remember how to do this. I'm I suppose to multiply?

OpenStudy (phi):

The only reason people study math is to give them an excuse to think (or puzzle out a problem) Say you were in a canoe but did not paddle, and drifted downstream. According to the info in this problem, how fast will you be moving ?

OpenStudy (freemap):

Umm I really don't mean to sound stupid, but 1 maybe. not very fast

OpenStudy (phi):

you will be going at a speed of "y" (they tell us let y represent the speed of the water.)

OpenStudy (phi):

though y is a letter, it means "speed of the current" (it's short-hand)

OpenStudy (phi):

If you start paddling you will go faster. You would add on the speed you can paddle they say Let x represent the speed of the boat (when paddling in still water) do you know how to show x added to y ?

OpenStudy (freemap):

hmm y+ x+y?

OpenStudy (phi):

just x+y that is how fast you go if you go downstream we might want to put parens around it , so we remember it as one thing: speed of the boat: (x+y) now we use speed * time = distance (x+y)* time = distance they don't tell us the distance, we let's call the distance "d" but we know the time. can you fill in the time with a number and write the full equation?

OpenStudy (freemap):

(x+y)*1=d

OpenStudy (phi):

looks good, except why are you using 1 for the time ? what does the question say the time is for going downstream ?

OpenStudy (freemap):

(x+y)*3.6h=d

OpenStudy (phi):

Write an expression for the distance traveled downstream using 3 h for the time.

OpenStudy (freemap):

(x+y)*3h=d

OpenStudy (phi):

ok. I think the h mean hours. Probably we should leave it off. so (x+y)*3 = d or 3(x+y)= d is the answer to the first part.

OpenStudy (phi):

now you need to the speed going upstream any ideas ?

OpenStudy (freemap):

(x+y)*3.6=d or 3.6(x+y)=d

OpenStudy (phi):

(x+y) is how fast you go downstream when you go upstream, you go slower

OpenStudy (freemap):

would it be x-y

OpenStudy (phi):

the river is taking you backwards at a speed of y , as you paddle upriver at a speed of x yes (x-y) is the speed going upstream

OpenStudy (freemap):

ok would we still multiply the time like (x-y)*3.6=d

OpenStudy (phi):

yes, exactly. now you have part a) 3(x+y)= d 3.6(x-y) = d

OpenStudy (phi):

now ***Set the expressions in Part (a) equal to each other. Then solve the equation for y. Show your work. *** we see that 3(x+y) = d and 3.6(x-y) also equals d. if both are equal to d, we can say 3(x+y)= 3.6(x-y) Does that make sense ?

OpenStudy (freemap):

It does make since d is the outcome in both equations

OpenStudy (freemap):

expressions i mean

OpenStudy (phi):

to solve, distribute the 3 on the left side. that means multiply 3 times x and times y ditto for the 3.6 on the other side

OpenStudy (freemap):

ok 3x+3y=3.6x-3.6y do we then add like terms?

OpenStudy (phi):

yes,

OpenStudy (phi):

3x+3y=3.6x-3.6y I would add 3.6y to both sides 3x + 3y + 3.6y = 3.6x -3.6y +3.6y

OpenStudy (phi):

you get 3x+6.6y = 3.6x now add -3x to both sides

OpenStudy (freemap):

6.6y=-0.833

OpenStudy (phi):

you lost the x? what is 3.6x - 3x ?

OpenStudy (freemap):

0.6x

OpenStudy (phi):

yes, so you get 3x+6.6y = 3.6x 6.6y = 0.6 x now divide both sides by 6.6 what do we get ?

OpenStudy (freemap):

11

OpenStudy (phi):

you get \[ y= \frac{0.6}{6.6} x\] that simplifies to \[ y = \frac{1}{11} x \]

OpenStudy (freemap):

Ok I get it

OpenStudy (phi):

**(c) What percent of the boat’s speed is the water current? *** I would find the ratio of y/x and change it to a percent. of course, we can't use just "y", but we know y is the same as x/11 so use x/11 instead. can you do that ?

OpenStudy (phi):

you do \[ \frac{y}{x}= y \cdot \frac{1}{x} \] but y is x/11 so \[ \frac{y}{x}= \frac{x}{11} \cdot \frac{1}{x} \]

OpenStudy (freemap):

ok, so if i multiply it turns back into 1x and 11x so I'm really sure how to solve this

OpenStudy (phi):

you should learn that if you have the same thing "up top" and "below" , they cancel when you multiply fractions, you multiply top times top and bottom times bottom \[ \frac{x \cdot 1 }{11 \cdot x}\] as you know we can change the order of the multiply (right ?) so it's the same as \[ \frac{1 \cdot x }{11 \cdot x}\] but that is the same as multiplying the two fractions \[ \frac{1 \cdot x }{11 \cdot x} = \frac{1}{11} \cdot \frac{x}{x}\]

OpenStudy (phi):

the last step, we "undid" the multiply the point is, we have x/x and anything divided by itself is 1 in other words we get \[ \frac{1}{11} \cdot 1 = \frac{1}{11} \]

OpenStudy (phi):

that is the long way. the short way is to say "x up top, x down below" cross off both \[ \frac{y}{x}= \frac{\cancel{x}}{11} \cdot \frac{1}{\cancel{x}} = \frac{1}{11}\]

OpenStudy (freemap):

Thats what I meant to say 1/11 because I did cross multiply

OpenStudy (phi):

no, you should not cross multiply. Try to follow what I posted up above. meanwhile, to answer the question, change 1/11 to a decimal, and then multiply by 100 to make it a percent.

OpenStudy (freemap):

ok, 0.11

OpenStudy (phi):

0.11 means 11/100 that is different from 1/11 to change it to a decimal, divide 11 into 1 (a calculator will do that) or type 1/11= into google

OpenStudy (freemap):

9.09 after multiplying 100

OpenStudy (phi):

and add a % sign (which is how we show we multiplied by 100) 9.09%

OpenStudy (freemap):

9.09% got it

OpenStudy (freemap):

Thank you so very much. I really appreciate your help.